OpenStudy (baldymcgee6):

In the course of an experiment you are measuring a quantity that is expected to change according to an exponential model. Your first measurement of the quantity gives a reading of 50 units, while three hours later the reading is 48. Construct the formula for the function that represents this quantity as a function of time.

OpenStudy (anonymous):

evidently it is decreasing at a rate of 4% every 3 hours, so a snap answer would be \(50\times (.94)^{\frac{t}{3}}\)

OpenStudy (anonymous):

you are probably supposed to find a model that looks like \[p(t)=p_0e^{kt}\] which means solving an exponential equation we can do that too if you like, but the first answer is much snappier

OpenStudy (baldymcgee6):

i think we're supposed to represent it as a decay formula, so like the p(t) function you posted

OpenStudy (anonymous):

ok then you want \[p(t)=p_0e^{kt}\] we start with \(p_0=50\) because that is how much you have at the beginning, so what we need is \(k\)

OpenStudy (anonymous):

you know when \(t=3\) you get 48, so you can write \[48=50e^{k\times 3}\] and solve for \(k\)

OpenStudy (baldymcgee6):

how do we solve for k? logarithms?

OpenStudy (anonymous):

\[48=50e^{3k}\] divide by 50 get \[.96=e^{3k}\] then yes, you need to take the log

OpenStudy (anonymous):

you get \[\ln(.96)=3k\] and so \[k=\frac{\ln(.96)}{3}\] whatever that is (need a calculator) then you are done

OpenStudy (baldymcgee6):

okay, thanks so much